OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (anonymous):

y -y1 = m(x -x1) y - (-1) = 4(x - (-3))

OpenStudy (anonymous):

y=4x+11 Beacuse -1=4(-3)+11 -1=-12+11 -1=-1

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

you are welcome! if you need anything else, just send it through here!